A German worker takes 400 hours to produce a car and 2 hours to produce a case of wine.A French worker takes 600 hours to produce a car and X hours to produce a case of wine.


a.for what value of x will gains from trade be possible ? Explain.


b . For what values of X will Germany export cars and import wine? Explain.

1
Expert's answer
2022-02-24T11:42:26-0500

Germany:

1 car = 200 cases of wine,

1 case of wine = 1/200 car

France:

1 car = 600/X cases of wine,

1 case of wine = x/600 car


(a) Gains from trade will be possible when X does not equal 3. When X=3, the opportunity cost of 1 car in France is 200 cases of wine, neither country has a comparative advantage. At all other values of X, a comparative advantage will exist.

(b) Germany will export cars and import wine for all values of X<3. When X<3, Germany has comparative advantage in producing car, and France has comparative advantage in producing wine. Therefore, Germany exports cars and imports wine